OpenStudy (nuccioreggie):

Simplify (4.5)(5)(−2). 45 4.5 −4.5 −45

OpenStudy (inowalst):

Multiply each number.

OpenStudy (anonymous):

-45

OpenStudy (anonymous):

fan and medal and @ me for more help @nuccioreggie

OpenStudy (welshfella):

multiply from left to right 4.5 * 5 = ? then multiply what you get by -2 Don't forget: a positive number * a negative gives a negative answer

OpenStudy (inowalst):

\[4.5 \times5\times-2\]\[4.5 \times5 = 22.5 \]\[22.5 \times-2= ?\]

OpenStudy (nuccioreggie):

I GOT d

OpenStudy (inowalst):

D is correct.

OpenStudy (inowalst):

\[22.5 \times -2=-45\]